# DrVoice: Parallel Speech-Text Voice Conversation Model via Dual-Resolution Speech Representations

### v2

@@ Abstract
Recent studies on end-to-end speech generation with large language models (LLMs) have attracted significant community attention, with multiple works extending text-based LLMs to generate discrete speech tokens. Existing approaches primarily fall into two categories: (1) Methods that generate discrete speech tokens independently without incorporating them into the LLM’s autoregressive process, resulting in text generation being unaware of concurrent speech synthesis. (2) Models that generate interleaved or parallel speech-text tokens through joint autoregressive modeling, enabling mutual modality awareness during generation. This paper presents DrVoice, a parallel speech-text voice conversation model based on joint autoregressive modeling, featuring dual-resolution speech representations. Whereas current methods utilize mainly 12.5Hz input audio representation, our proposed dual-resolution mechanism reduces the input frequency for the LLM to 5Hz. Experimental results on Spoken Question Answering benchmarks demonstrate that DrVoice establishes new state-of-the-art (SOTA) performance among similar size speech foundation models with relative small amount of data.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) (OpenAI, 2023; Yang et al., 2024) have demonstrated remarkable performance across various NLP tasks, sparking growing interest in extending their capabilities to multimodal domains encompassing vision and audio. Developments in spoken dialogue systems are critical to human-computer interaction, as natural human communication inherently relies on verbal exchanges. The release of GPT-4o (OpenAI, 2024b) is a significant milestone in this field, showing the potential of LLM-based spoken dialogue systems for seamless and natural interactions with users.
@@
LLM-based spoken dialogue systems can be generally categorized into cascaded systems and end-to-end (E2E) systems, with the distinction lying in whether the backbone LLM can directly comprehend speech representations and generate speech outputs. Early cascaded systems integrate separately trained Automatic Speech Recognition (ASR), LLM, and Text-to-Speech (TTS) modules. Such systems inherently suffer from error accumulation, loss of acoustic details (e.g., tone, emotion), and significant latency. Alternatively, the ASR module could be eliminated by introducing audio understanding foundation models (Chu et al., 2023, 2024). E2E systems have emerged to further eliminate ASR and TTS modules, and establish direct connections between speech representations and LLMs. However, training LLMs to generate highly intelligent speech outputs remains challenging. E2E systems struggle with the quality of speech token generation, primarily due to inefficient utilization of textual information during speech token generation.
@@
Recent research efforts on E2E models to address the above challenges have focused on two primary directions Chen et al. (2024): Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models. Text-Driven Speech Models refer to systems in which LLMs process speech representations as input, produce textual responses, and utilize the LLM’s representations as input to a speech decoder for speech generation. In contrast, Joint Speech-Text Models involve LLMs taking speech representations as input and generating both text tokens and speech tokens simultaneously. The key distinction lies in whether the generation of speech tokens can influence the subsequent generation of text tokens within the LLM. This feedback loop of speech tokens to LLM is present in Joint models but absent in Text-Driven models. Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024; KimiTeam et al., 2025). Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM. While both methodologies explicitly incorporate textual guidance into speech token generation to leverage the LLM’s capabilities, they have distinct limitations. The architecture of Text-Driven Speech Models creates a unidirectional information flow where text generation is completed before speech synthesis begins. This prevents the LLM from conditioning its textual output on the generated acoustic tokens, thus limiting its ability to control fine-grained paralinguistic attributes like emotion and prosody.
@@
Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge. Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials Chen et al. (2024); hence, in this work, we focus on enhancing joint speech-text modeling. Recently, Kimi-Audio KimiTeam et al. (2025) has set a new state-of-the-art (SOTA) in joint speech-text modeling. However, this approach faces notable challenges. It not only requires extensive data for training but also incurs significant computational costs due to its 12.5Hz audio representation. Furthermore, this high token rate creates a frequency mismatch with the much lower rate of text tokens ($\sim$3Hz) (Chen et al., 2024), which can dilute semantic information and consequently hinder the full utilization of the LLM’s core capabilities.
@@
In this work, we propose DrVoice, a novel parallel speech-text voice conversation model with Dual-Resolution Speech Representations (DRSR), which compresses the input down to 5Hz. DRSR is proposed mainly to alleviate the temporal resolution discrepancy between speech and text tokens. For speech comprehension, we introduce a grouping mechanism that maps 25Hz discrete audio tokens to 5Hz speech representations, which better match the pace of text. During generation, combined speech-text embeddings form the assistant’s autoregressive input. The hidden states captured from shared LLM layers are then passed in parallel to a text head for text token prediction, and a Speech Refined Head (SRH) to generate the corresponding ungrouped speech tokens. Due to input-output resolution discrepancies, each LLM inference triggers $k$ parallel SRH executions, where $k$ is the grouping factor. To further enhance the model’s reasoning and the coherence of its output, we incorporate a Chain-of-Modality (CoM) (Zhang et al., 2023a) strategy. This approach prompts the model to first generate a complete response in text, allowing it to structure its thoughts before engaging in parallel speech-text generation. This intermediate reasoning step serves to improve modality alignment and reduce errors. For practical applications, we design system prompts to control the output modes, enabling text-only output, direct parallel speech-text output, or the CoM-enhanced parallel output. Experimental results demonstrate the efficacy of CoM in significantly improving performance. For reproducibility, we will release all source code and model checkpoints.
@@
Our contributions are three-fold: 1) We propose DrVoice, a novel parallel speech-text conversation model featuring Dual-Resolution Speech Representations (DRSR). This core architectural innovation effectively alleviates the temporal resolution mismatch between speech and text tokens, reducing computational costs and better preserving the LLM’s semantic processing capabilities. 2) We introduce a CoM-Mixing training strategy which acts as a curriculum, using the structured reasoning of CoM to scaffold directly speech generation task. 3) Through comprehensive experiments, we demonstrate that DrVoice achieves new state-of-the-art (SOTA) performance on multiple spoken question answering benchmarks, validating the effectiveness of our proposed architecture and strategies.
@@ 2 Related Work
Two primary directions exist for converting continuous speech signal into processable sequences: continuous representations, e.g., Whisper (Radford et al., 2022), and discrete representations. Since LLMs generate discrete tokens, continuous representations face challenges in direct integration with LLMs for speech generation. In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (Défossez et al., 2023), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (Hassid et al., 2023) or supervised learning as in CosyVoice (Du et al., 2024a, b), and prioritized for linguistic content (Hsu et al., 2021). While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (Zhang et al., 2023b), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (Zeng et al., 2025; Zhang et al., 2024b, 2023a; Borsos et al., 2023). In this work, we select HuBERT-derived semantic tokens (Hassid et al., 2023) over other paradigms due to their competitive semantic capabilities and compatibility with a more lightweight, diffusion-free vocoder for direct speech synthesis.
E2E Speech Foundation Models.
@@
Existing E2E speech foundation models can be categorized into Text-Driven Speech Models and Joint Speech-Text Models. Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adaptor, LLM, and a streaming speech decoder, and can simultaneously generate text and speech. Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker specializing in speech token production. Since Thinker cannot receive speech tokens during generation, the framework inherently limits awareness of speech token generation states, constraining applications such as full-duplex voice conversation.
@@
Joint Speech-Text Models explore two different architectures, Interleaved and Parallel. Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025; Long et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024; KimiTeam et al., 2025) conduct parallel decoding. The interleaved model VITA-Audio Long et al. (2025) achieves efficient audio generation by introducing multiple cross-modal token prediction (MCTP) modules. It demonstrates superior performance across multiple benchmarks while maintaining high-quality speech synthesis. The parallel model Moshi (Défossez et al., 2024) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations. To mitigate limitations of discrete tokens in audio understanding, Kimi-Audio (KimiTeam et al., 2025) introduce dual-tokenizer combining discrete semantic tokens with continuous Whisper Radford et al. (2023) features, preserving both semantic and acoustic information. In order to further enhance the efficiency of both training and inference and alleviate the issue of granularity misalignment between the text and speech modalities, our approach leverages DRSR to reduce the LLM input frame rate to 5Hz, without sacrificing performance.
@@ 3 Methodology
Figure 1 illustrates the architecture of DrVoice. The system consists of three main components: (1) Speech Tokenizer and Speech Encoder are designed to process the wave of speech into hidden representation, (2) a Multimodal Large Language Model (MLLM) consists of shared LLM layer, text head, and speech refined head (SRH) for token generation, and (3) Speech Detokenizer to generate wave from speech tokens. The system operates through multimodal input encoding and coordinated speech-text output generation. During inference, user inputs (text or speech) are first mapped to a unified semantic space, processed by MLLM to produce parallel speech-text responses through SRH and text head. To effectively train this system, we introduce the CoM-Mixing Training strategy.
@@ 3.1 Speech Tokenization and Detokenization
Semantic tokens have been widely used for speech tokenization (Zhang et al., 2023a; Borsos et al., 2023), due to their strong alignment with text (Zhang et al., 2023b); therefore, we employ HuBERT (Hsu et al., 2021) as the Speech Tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$, where $T$ is the speech token sequence length. For speech detokenization, HiFi-GAN (Kong et al., 2020) is selected to regenerate speech waveform from speech tokens $\mathbf{S}$, since it is a non-autoregressive neural vocoder that efficiently generates high-fidelity waveforms and reduces the overall latency. Inspired by Kimi-Audio KimiTeam et al. (2025) and aiming for improved audio understanding, we utilize the Whisper-Large-v3 Speech Encoder to extract continuous audio representations. Subsequently, an Adapter is introduced to downsample the temporal resolution of these representations and align their hidden dimension with that of the LLM. Finally, the processed continuous representations are stacked with discrete ones and fed into the LLM.
@@
Figure 1: Overview of DrVoice. User speech inputs are tokenized, grouped, and encoded by the MLLM for autoregressive text and speech token prediction. The generated speech tokens are then converted to speech waveform by the speech detokenizer. Note that SRH through $k$ autoregressive forward passes to generate $k$ speech tokens, where $k$ is the grouping factor.

### v3

@@ Abstract
Recent studies on end-to-end (E2E) speech generation with large language models (LLMs) have attracted significant community attention, with multiple works extending text-based LLMs to generate discrete speech tokens. Existing E2E approaches primarily fall into two categories: (1) Methods that generate discrete speech tokens independently without incorporating them into the LLM’s autoregressive process, resulting in text generation being unaware of concurrent speech synthesis. (2) Models that generate interleaved or parallel speech-text tokens through joint autoregressive modeling, enabling mutual modality awareness during generation. This paper presents DrVoice, a parallel speech-text voice conversation model based on joint autoregressive modeling, featuring dual-resolution speech representations. Notably, while current methods utilize mainly 12.5Hz input audio representation, our proposed dual-resolution mechanism reduces the input frequency for the LLM to 5Hz, significantly reducing computational cost and alleviating the frequency discrepancy between speech and text tokens and in turn better exploiting LLMs’ capabilities. Experimental results demonstrate that DrVoice-7B establishes new state-of-the-art (SOTA) on OpenAudioBench and Big Bench Audio benchmarks, while achieving performance comparable to the SOTA on VoiceBench and UltraEval-Audio benchmarks, making it a leading open-source speech foundation model in $\sim$7B models.
@@ 1 Introduction
Developments in spoken dialogue systems are critical to human-computer interaction, as natural human communication inherently relies on verbal exchanges. Recently, Large Language Model (LLM) based spoken dialogue systems, exemplified by systems like GPT-4o (OpenAI, 2024b), demonstrate great potential for seamless and natural interactions with users. LLM-based spoken dialogue systems can be generally categorized into cascaded systems and end-to-end (E2E) systems, with the distinction lying in whether the backbone LLM can directly comprehend speech representations and generate speech outputs. Early cascaded systems integrate separately trained Automatic Speech Recognition (ASR), LLM, and Text-to-Speech (TTS) modules. Such systems inherently suffer from error accumulation, loss of acoustic details (e.g., emotion), and significant latency. Alternatively, the ASR module could be eliminated by introducing audio understanding foundation models (Chu et al., 2023, 2024). E2E systems have emerged to further eliminate ASR and TTS modules, and establish direct connections between speech representations and LLMs. However, training LLMs to generate highly intelligent speech outputs remains challenging. E2E systems struggle with the quality of speech token generation, primarily due to inefficient utilization of textual information during speech token generation.
@@
Recent works on E2E models to address these challenges have focused on two primary directions (Chen et al., 2024a): Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models. Text-Driven Speech Models refer to systems in which LLMs process speech representations as input, produce textual responses, and utilize the LLM’s representations as input to a speech decoder for speech generation. In contrast, Joint Speech-Text Models involve LLMs taking speech representations as input and generating both text tokens and speech tokens simultaneously. The key distinction lies in whether speech token generation can influence the subsequent generation of text tokens within the LLM: this feedback loop of speech tokens to LLM is present in Joint models but absent in Text-Driven models. Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024a; KimiTeam et al., 2025). Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM.
@@
While both Text-Driven Speech Models and Joint Speech-Text Models explicitly incorporate textual guidance into speech token generation to leverage the LLM’s capabilities, they have distinct limitations. The architecture of Text-Driven Speech Models creates a unidirectional information flow where text generation is completed before speech synthesis begins. This prevents the LLM from conditioning its textual output on the generated speech tokens, thus limiting its ability to explore fine-grained paralinguistic attributes like emotion and prosody. On the other hand, Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge. Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials (Chen et al., 2024a); hence, in this work, we focus on enhancing joint speech-text modeling. Recently, Kimi-Audio (KimiTeam et al., 2025) sets a new state-of-the-art (SOTA) in joint speech-text modeling. However, this approach still suffers from notable limitations. It not only requires extensive training data but also incurs significant computational costs due to its 12.5Hz audio representation. Furthermore, the high token rate creates a frequency mismatch with the much lower rate of text tokens ($\sim$3Hz) (Chen et al., 2024a), which can dilute semantic information and consequently hinder the full utilization of the LLM’s core capabilities.
@@
In this work, we propose DrVoice, a novel parallel speech-text voice conversation model with Dual-Resolution Speech Representations (DRSR). For speech comprehension, we introduce a grouping mechanism that maps 25Hz discrete audio tokens to 5Hz speech representations, effectively alleviating the temporal resolution discrepancy between speech and text tokens. During generation, combined speech-text embeddings form the assistant’s autoregressive input. The hidden states captured from shared LLM layer are then passed in parallel to a Text Head for text token prediction and a Speech Refined Head (SRH) to generate the corresponding ungrouped speech tokens.
To further enhance the model’s reasoning and coherence of its output, we incorporate a Chain-of-Modality (CoM) (Zhang et al., 2023a) strategy. CoM prompts the model to first generate a complete response in text, allowing it to structure its thoughts before engaging in parallel speech-text generation. This intermediate reasoning step improves modality alignment and reduces errors. We design system prompts to control the output modes, enabling text-only output, direct parallel speech-text output, or the CoM-enhanced parallel output. We also develop a Core-Cocktail training strategy to refine model optimization and LLM knowledge retention.
@@
Our contributions are three-fold: 1) We propose DrVoice, a novel parallel speech-text conversation model featuring Dual-Resolution Speech Representations (DRSR). This core architectural innovation effectively alleviates the temporal resolution mismatch between speech and text tokens, reducing computational costs and better preserving the LLM’s semantic processing capabilities. 2) We introduce two new training strategy, including a CoM-Mixing training strategy acting as a curriculum, using the structured CoM reasoning to scaffold speech generation, and a Core-Cocktail training strategy for retaining the knowledge of LLMs. 3) Our comprehensive experimental results reveal that DrVoice-7B achieves new SOTA performance on prominent benchmarks like OpenAudioBench (audio understanding) and Big Bench Audio (reasoning and understanding capabilities of audio-processing models), alongside performance on par with the SOTA on other benchmarks such as VoiceBench (benchmarking LLM-based voice assistants) and UltraEval-Audio (both speech understanding and generation), solidifying its position as a premier open-source speech foundation model among $\sim$7B models.
@@ 2 Related Work
Speech Tokenization. Two primary directions exist for converting continuous speech signal into processable sequences: continuous representations, e.g., Whisper (Radford et al., 2022), and discrete representations. Since LLMs generate discrete tokens, continuous representations face challenges in direct integration with LLMs for speech generation. In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (Défossez et al., 2023), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (Hassid et al., 2023) or supervised learning as S3Tokenizer using in CosyVoice (Du et al., 2024a, b, 2025), and prioritized for linguistic content (Hsu et al., 2021). While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (Zhang et al., 2023b), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (Zeng et al., 2025; Zhang et al., 2024b, 2023a; Borsos et al., 2023). In this work, we use S3Tokenizer as the speech tokenizer due to its solid semantic capabilities and compatibility with the strong Speech Detokenizer from CosyVoice for synthesis. S3Tokenizer is a supervised semantic speech tokenizer based on the pre-trained SenseVoice-Large model (An et al., 2024) that enhances the semantic relationship of extracted tokens. S3Tokenizer is robust to data noise, and reduces the reliance on clean data.
@@
E2E Speech Foundation Models. Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adapter, LLM, and a streaming speech decoder, and can simultaneously generate text and speech. Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker handling speech token production. Since Thinker cannot receive speech tokens during generation, the framework inherently limits awareness of speech token generation states, constraining applications such as full-duplex voice conversation.
@@
Joint Speech-Text Models explore two different architectures, Interleaved and Parallel. Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025; Long et al., 2025; Wu et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024a; KimiTeam et al., 2025) conduct parallel decoding. The parallel model Moshi (Défossez et al., 2024) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations. To mitigate limitations of discrete tokens in audio understanding, Kimi-Audio (KimiTeam et al., 2025) introduces dual-tokenizer combining discrete semantic tokens with continuous Whisper (Radford et al., 2023) features, preserving both semantic and acoustic information. In order to further enhance the efficiency of both training and inference and alleviate the issue of granularity misalignment between the text and speech modalities, our approach leverages DRSR to reduce the LLM input frame rate to 5Hz, without sacrificing performance.
@@ 3 Methodology
Figure 1 illustrates the architecture of DrVoice. The system consists of three main components: (1) Speech Encoder and Speech Tokenizer process the speech wave into hidden representations for the user end and the assistant end respectively, (2) a Multimodal Large Language Model (MLLM) consists of shared LLM layer, a Text Head, and a Speech Refined Head (SRH) for token generation, and (3) Speech Detokenizer to generate wave from speech tokens. The system operates through multimodal input encoding and coordinated speech-text output generation. During inference, user inputs (text or speech) are first mapped to a unified semantic space, processed by MLLM to produce parallel speech-text responses through SRH and text head. To effectively train this system, we introduce the CoM-Mixing Training and Core-Cocktail Training strategies.
@@ 3.1 Speech Tokenization and Detokenization
Aiming for enhanced audio understanding, we utilize Whisper-Large-v3 Speech Encoder to extract continuous audio representations at the user end. Subsequently, an Adapter is introduced to downsample the temporal resolution of these representations and align their hidden dimension with that of the LLM. Semantic tokens have been widely used for speech tokenization (Zhang et al., 2023a; Borsos et al., 2023), due to their strong alignment with text (Zhang et al., 2023b); therefore, we employ S3Tokenizer (Du et al., 2024a, b, 2025) as the Speech Tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$ at the assistant end, where $T$ is the speech token sequence length. For speech detokenization, conditioned on speaker embeddings capturing acoustic details such as timbre, the Flow Matching model (Lipman et al., 2023) converts speech tokens $\mathbf{S}$ into the Mel spectrum for a given speaker. Finally, a pre-trained vocoder HiFi-GAN (Kong et al., 2020) transforms the Mel spectrum back into audio signal.
@@
Figure 1: Overview of DrVoice. User speech inputs are tokenized, grouped, and encoded by the MLLM for autoregressive text and speech token prediction. The MLLM consists of Shared LLM Layer, a Text Head, and a Speech Refined Head (SRH) for token generation. The generated speech tokens are then converted to speech waveform by the speech detokenizer. Note that SRH generates $k$ speech tokens through $k$ autoregressive forward passes, where $k$ is the grouping factor.
3.2 Multimodal Large Language Model (MLLM)
Parallel Joint Speech-Text Model. Inspired by Moshi (Défossez et al., 2024), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold. We focus on performing modality alignment exclusively at the assistant end. This design adheres to the asymmetric nature of human-machine interactions: while user inputs are typically unimodal (either text or speech), the assistant’s responses can be a coordinated multimodal output.
$c_{t}=E_{\text{speech}}(s_{t})+E_{\text{text}}(t_{t})$ (1)
where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively. Since the lengths of speech tokens and text tokens are typically mismatched, the shorter sequence is padded with a special token <|SIL|>to align them for each utterance. The autoregressive generation process is as follows:
$P(y_{t}|y_{<t},x)=\prod_{i=1}^{t}P(y_{i}|y_{<i},x)$ (2)
where $x$ is the input sequence and $y_{t}=(s_{t},t_{t})$ denotes the joint speech-text output at timestep $t$. This unified formulation enables seamless integration of speech and text generation within a single autoregressive framework. To preserve the intrinsic linguistic understanding and generation capabilities of pretrained text-LLMs while enabling cross-modal interactions, three key methodological designs are introduced for intermodal coordination, including Speech Token Grouping and Ungrouping, and Speech Refined Head.
Speech Token Grouping. A crucial parameter for discrete tokens is the sampling rate, which determines the input/output sequence length. GLM-4-Voice (Zeng et al., 2025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz. Hence, our work adopts 25Hz sampling rate. To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (Chen et al., 2024a), a grouping mechanism is designed:
where $\mathbf{s}_{j}$ denotes speech tokens, $\|$ represents feature concatenation, and $k$ is the grouping factor determined by the ratio between speech token rates and text token rates. This design compresses the speech token length from $T$ to $T/k$ and the resulting grouped speech representations align better with text. Notably, different from Chen et al. (2024a) that employs linear projection of audio logits into group-sized representations for parallel multi-token prediction, DrVoice strategically designs an ungrouping mechanism and incorporates a dedicated Speech Refined Head (SRH) to enable autoregressive generation of individual speech tokens.
Speech Refined Head (SRH). SRH is proposed to enhance speech generation capabilities. It utilizes the last hidden state of the shared LLM layer (SLLM) as conditional input and incorporates contextual speech information to autoregressively generate speech tokens. While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024a; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details. To address this limitation, DrVoice performs an ungrouping process as follows: the SLLM’s final hidden state is mapped to group-sized embeddings via linear projection
and time splitting
where $\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k}$. The resulting $\mathbf{H}$ serves as the conditional input for SRH that autoregressively generates speech tokens. Following our practice on SLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated. Speech token prediction is trained to maximize the conditional probability:
The E2E training objective $\mathcal{L}_{\text{MLLM}}$ integrates the two losses through multi-task learning:
$\mathcal{L}_{\text{TH}}=-\sum_{i=1}^{T}\log P(t_{i}|c_{<i},\mathbf{g}),$ (8)
where $\mathcal{L}_{\text{TH}}$ is the autoregressive loss over the text head, $\lambda$ and $\mu$ are hyperparameters.
3.3 Training Strategy
Initialization. The Speech Encoder is initialized with the weights of Whisper-Large-v3, while the Shared LLM Layer is initialized using Qwen2.5. Additionally, the pre-trained Speech Tokenizer and Detokenizer from CosyVoice are employed and kept frozen throughout the entire training process. We initialize SRH with a pre-trained TTS model.
CoM-Mixing Training. The chain-of-modality (CoM) methodology (Zhang et al., 2023a) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical. Furthermore, practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs. Seven interaction patterns for diverse input-output requirements are summarized in Table 1. System prompts are employed to regulate the model’s output behavior. For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs. Detailed prompts are shown in Appendix B. During inference, specifying these system prompts enables the generation of desired output results. For CoM-Mixing training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model. During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement. This flexibility ensures adaptability to varying modality generation demands.